# Beyond the Encoder: Joint Encoder-Decoder Contrastive Pre-Training Improves Dense Prediction

### v2

@@ Abstract
Contrastive learning methods in self-supervised settings have primarily focused on pre-training encoders, while decoders are typically introduced and trained separately for downstream dense prediction tasks. However, this conventional approach overlooks the potential benefits of jointly pre-training both encoder and decoder. In this paper, we propose DeCon, an efficient encoder-decoder self-supervised learning (SSL) framework that supports joint contrastive pre-training. We first extend existing SSL architectures to accommodate diverse decoders and their corresponding contrastive losses. Then, we introduce a weighted encoder-decoder contrastive loss with non-competing objectives to enable the joint pre-training of encoder-decoder architectures. By adapting an established contrastive SSL framework for dense prediction tasks, DeCon achieves new state-of-the-art results: on COCO object detection and instance segmentation when pre-trained on COCO dataset; across almost all dense downstream benchmark tasks when pre-trained on COCO+ and ImageNet-1K. Our results demonstrate that joint pre-training enhances the representation power of the encoder and improves performance in dense prediction tasks. This gain persists across heterogeneous decoder architectures, various encoder architectures, and in out-of-domain limited-data scenarios.
**footnotetext: Equal contributions for this work. Co-first authors are ordered in alphabetical order.
@@ 1 Introduction
The growing demand for efficient Deep Learning (DL) methods stems from their ability to address complex tasks. While traditional approaches like supervised learning can yield high-performing models, they depend on large volumes of high-quality annotated data. Obtaining such annotations is often labor-intensive, expensive, and sometimes impractical, even when abundant data is available. A common workaround is to use pre-trained models, typically encoders trained from scratch on large-scale datasets, to boost performance on downstream tasks with limited labels. Typically, an encoder is pre-trained to capture a representation of the input data, which can then be fine-tuned for downstream tasks such as classification or segmentation. During the downstream fine-tuning phase, randomly initialized layers are typically added on top of the encoder, and the complete architecture is retrained using a supervised learning approach on an annotated dataset. ImageNet [15] is a widely utilized labeled dataset for such pre-training. However, models pre-trained on ImageNet classification task often transfer poorly to downstream dense prediction tasks, such as detection and segmentation [56, 47, 41].
@@
In this work, we propose a Decoder-aware contrastive learning (DeCon) approach, an encoder-decoder framework that jointly pre-trains encoder and decoder in an SSL manner. DeCon  enhances the representation power of the encoder and prepares it more effectively for integration with decoders in downstream dense prediction tasks. The contributions of this paper are as follows:
  • We propose DeCon-SL, a single-level joint encoder-decoder adaptation of contrastive dense SSL frameworks. We show that jointly pre-training the encoder and decoder enhances the representation power of the encoder and improves downstream dense prediction performance, even when transferring only the pre-trained decoder.
  • We further extend this adaptation by introducing DeCon-ML, which builds on DeCon-SL with a multi-level decoder loss function and an encoder-level channel dropout to promote a comprehensive feature utilization of the encoder. DeCon-ML achieves new state-of-the-art results in COCO object detection and instance segmentation when pre-trained on COCO dataset; and across almost all dense downstream benchmark tasks when pre-trained on COCO+ and ImageNet-1K.
  • We demonstrate that DeCon improves the encoder performance in downstream tasks compared to encoder-only frameworks, even when the decoders used for pre-training and fine-tuning differ.
  • We show that DeCon can achieve gains without increasing parameter count or GPU requirements relative to the original framework.
  • We demonstrate that DeCon remains effective in out-of-domain downstream tasks and yield even greater benefits in scenarios with limited annotated data for fine-tuning.
2 Related Works
@@
Recent SSL methods prioritize generalizability and move away from hand-crafted pretext tasks. These approaches can be broadly categorized into generative and contrastive methods. Generative approaches aim to learn meaningful representations by modeling the underlying data distribution. This is typically achieved through reconstruction-based methods, where the model predicts corrupted parts of the input [23, 1, 17, 53], or through adversarial methods, where a generator model learns to synthesize realistic samples to challenge a discriminator, thereby encouraging the encoder to extract informative semantic features [50, 26]. Contrastive methods, which are the focus of this research, learn data representations by maximizing the similarity between features extracted from two augmented views of the same image. To avoid model collapse—a phenomenon where the model maps all inputs to the same representation—strategies such as negative sampling [22, 8], stop-gradient mechanisms [9], variance regularization [2, 3], momentum encoders [22], and other architectural and optimization techniques have been employed [5].
@@
Encoder-only versus encoder-decoder pre-training learning: Encoder-decoder architectures with skip connections between the encoder and decoder layers (i.e., the U-Net family of models) have achieved state-of-the-art results in many segmentation tasks [43, 27, 55]. A common approach for pre-training such models in a self-supervised manner is image reconstruction [30, 4, 6]. However, the presence of skip connections allows a substantial amount of information to bypass the encoder’s final layer—often referred to as the bottleneck layer–resulting in lower-quality embeddings generated by the encoder. Our proposed approach preserves skip connections while addressing the issue of inferior representations generated by the decoder.
Encoder-only contrastive learning frameworks are built solely around an encoder, incorporating components such as predictors and projectors. However, they do not include a decoder. After pre-training, a randomly initialized decoder is attached to the pre-trained encoder in dense prediction downstream tasks to form a complete architecture. While some generative SSL frameworks use a decoder for pre-training [14, 40, 23], to the best of our knowledge, no work has explored pre-training vision models with both encoder and decoder loss in a contrastive learning framework to learn informative data representations.
@@ 3.1 DeCon-SL
Given a vision SSL framework that pre-trains an encoder using a teacher-student architecture [25], we retain the encoders and any existing “auxiliary layers” (e.g., predictors, projection heads), and add decoders and their corresponding auxiliary layers for both teacher and student networks. For all models, the auxiliary layers are designed to match those of the encoder, with adjustments applied to align with the decoder’s output feature size. We obtain two losses, one is computed from the encoder features while the other one is computed from the decoder features. Consequently, the loss function is defined as follows:
@@ 3.3 Implementation details
Architecture: We use a ResNet-50 encoder [21] as a backbone and implement a Fully Convolutional Network (FCN) [34] and a skip-connected Feature Pyramid Network (FPN) [32] as two decoder architectures in the DeCon framework. We follow the mmseg [11] FCN implementation used in SlotCon [48] and customize the FPN implementation from Detectron2 [49] version 0.6. Details on these decoder architectures are provided in Supplementary Material Section 1. We use DeCon to adapt the SlotCon [48] framework. Additional experiments are conducted using DenseCL [47] and PixPro [52] framework or a ConvNeXt-small [33] backbone.
@@
We adapt in a similar way DenseCL and PixPro into a DeCon-SL framework using an FCN decoder. More details are given in Supplementary Material 2. We adapt DenseCL implementation from mmselfsup package [12] version 1.0.0 and PixPro from the original implementation.
@@
Although the backbone pre-trained remains unchanged, adding multiple decoder levels and associated loss functions increases the overall number of parameters in the DeCon-ML architecture. To enable a fair comparison with the original SlotCon framework using a similar parameter budget, we implemented a reduced version of DeCon-ML that includes only the first two decoder levels originating from the encoder bottleneck. Additionally, the hidden dimension in the decoder projector was reduced from 4096 to 2048. This smaller version, which matches the parameter count of the original SlotCon framework (when $\alpha=0$), is referred to as DeCon-ML-S, while the full model with four decoder levels and unmodified configuration is denoted DeCon-ML-L. Detailed figures of DeCon-SL, DeCon-ML-S and DeCon-ML-L adaptations of SlotCon are available in the Supplementary Material Section 3.
@@
Pre-training setup: We pre-train DeCon on COCO 2017 dataset [31], COCO+ (train2017 + unlabeled2017 datasets) and ImageNet-1K datasets [15]. The augmentations and hyperparameters—learning rate, weight decay, schedulers, input shape, momentum, and so on—are kept the same as used in the original SSL framework. Models are pre-trained on COCO and COCO+ for 800 epochs, and for 200 epochs on ImageNet. The SlotCon ResNet-50 models are pre-trained using a LARS optimizer, with a total batch size of 512 across all available GPUs. The base learning rate is 1.0 scaled with the batch size and updated with a cosine learning rate decay schedule using weight decay of $10^{-5}$ and a warm-up period of 5 epochs. The ConvNeXt-small backbone is pre-trained in SlotCon and DeCon-ML using an AdamW optimizer [35], a base learning rate of 0.002 and weight-decay 0.05. Remaining hyperparameters are kept the same as for the ResNet-50 implementation. The pre-trainings using FPN decoders are done using eight NVIDIA H100 GPUs (80GB each). The pre-trainings without decoder and with FCN decoders are conducted on one 80GB NVIDIA A100 GPU.
For DenseCL, DeCon-SL is pre-trained using an SGD optimizer with a total batch size of 256. The base learning rate is 0.3 and is updated using a cosine learning rate decay schedule (0.0001 weight decay, 0.9 momentum). The pre-training experiments are conducted with eight NVIDIA A100 GPUs (40GB each). For PixPro, DeCon-SL is pre-trained using a LARS optimizer for 800 epochs with a batch size of 512 and a learning rate of 1.0 with a weight decay of 1e-5. The pre-trainings are conducted with eight NVIDIA A100 GPUs (40GB each).
To assess the DeCon framework on non scene-centric data, we pre-train SlotCon and a DeCon-SL adaptation of SlotCon with an FCN decoder on two small-scale datasets: ISIC 2017 and REFUGE described in Sec. 3.4. For ISIC dataset, we pre-train for 800 epochs with a batch size of 256 and a base learning rate of 1.0 linearly scaled with the batch size. For REFUGE, we pre-train for 2400 epochs with a batch size of 192 and the same learning rate as for ISIC. These pre-trainings are performed with one NVIDIA A6000 GPU with 48GB of memory.
@@ 3.4 Evaluation protocol
We evaluate the different pre-trained models on object detection, semantic segmentation, and instance segmentation as downstream tasks. Unless stated otherwise, we only transfer the pre-trained encoder for the fine-tuning experiments. We initialize the network with a pre-trained backbone and fine-tune end-to-end in all cases. All fine-tuning results reported in this paper are averaged over three independent runs. The pre-trainings, however, are only run once. All fine-tuning experiments related to the same dataset are run on a similar platform with the same hardware. We perform object detection and instance segmentation using Detectron2 package [49] version 0.6, whereas mmsegmentation [11] version 0.30.0 is used for semantic segmentation tasks.
Object detection and instance segmentation: We perform object detection and instance segmentation on COCO dataset [31]. We fine-tune a Mask R-CNN with FPN backbone following SlotCon [48].
Semantic segmentation: We evaluate the semantic segmentation performance of a pre-trained ResNet-50 encoder with an FCN decoder on Pascal VOC [18] and Cityscapes [13] following SlotCon [48] and PixCon [39].
Semantic segmentation on out-of-domain datasets: We assess generalization on REFUGE [38] and ISIC 2017 [10] datasets using the same ResNet-50-FCN architecture, fine-tuned on 5, 25, and 100% of the training data.
More details on the training setups and hyperparameters are provided in the Supplementary Material Section 2.
@@ 4.1 Experiments
Table 1: Performance of DeCon-SL ($\alpha=0.25$) and DeCon-ML Small and Large ($\alpha=0$ and $dropout=0.5$) adaptations of SlotCon trained with a ResNet-50 backbone. The architecture was pre-trained for 800 epochs on COCO and COCO+ and for 200 epochs on ImageNet-1K, and then fine-tuned on various segmentation and object detection tasks. Only the pre-trained encoder was transferred for the fine-tuning. Results for PixCon pre-trained on ImageNet-1K are not reported, as the corresponding checkpoint or evaluation was not made available in the original publication. All fine-tuning results were averaged over three runs. means that we reimplemented the framework, * means that the results are taken from the original paper.
Pret. Dataset Framework Pret. Dec. Object Detection Instance Segmentation Semantic Segmentation
COCO COCO VOC City
AP AP50 AP75 AP AP50 AP75 mIoU mIoU
COCO PixCon [39]* - 40.81 60.97 44.80 36.80 57.93 39.62 72.95 76.62
SlotCon [48] - 40.81 60.95 44.37 36.80 57.98 39.54 71.50 75.95
DeCon-SL FCN 40.97 61.22 44.81 36.92 58.12 39.78 73.01 76.21
DeCon-ML-S FPN 40.97 61.20 44.71 36.94 58.20 39.63 72.80 76.21
DeCon-ML-L FPN 41.18 61.38 44.91 37.12 58.35 39.94 72.92 76.45
COCO+ PixCon [39]* - 41.2 - - 37.1 - - 73.9 77.0
SlotCon [48] - 41.63 62.10 45.67 37.57 59.07 40.45 73.93 76.43
DeCon-SL FCN 41.86 62.43 45.73 37.75 59.40 40.48 74.46 76.65
DeCon-ML-L FPN 42.08 62.42 46.13 37.84 59.41 40.75 75.36 77.00
ImageNet-1K SlotCon [48] - 41.69 62.07 45.59 37.59 58.97 40.49 75.02 76.15
DeCon-ML-L FPN 41.80 62.12 45.73 37.73 59.08 40.68 75.40 76.51
ResNet-50 backbone: Table 1 presents the results of fine-tuning a ResNet-50 backbone pre-trained with the two best pre-training frameworks for dense downstream tasks and DeCon. A more comprehensive comparison with additional previously published methods is provided in the Supplementary Material 4. DeCon-ML-L is consistently establishing a new SATO for all tasks, except when pre-training is performed on COCO dataset for Pascal VOC and Cityscapes semantic segmentation fine-tuning, where it reduces the gap between PixCon and SlotCon. This supports the fact that the DeCon adpatation is robust across datasets. Secondly, we can see that all of DeCon versions consistently improve upon the base SlotCon framework for any of the pre-training datasets used. Notably, DeCon-ML-S, whose parameter count is similar to SlotCon, exhibits better performance than SlotCon.
Table 2: Performance of DeCon-ML-L ($\alpha=0$ and $dropout=0.5$) adaptation of SlotCon using a ConvNeXt-small encoder. The architecture was pre-trained for 800 epochs on COCO, and then fine-tuned on various segmentation and object detection tasks. All fine-tuning results were averaged over three runs.
@@
ConvNeXt backbone: Table 2 presents the results of the fine-tuning of models pre-trained with a ConvNeXt-small backbone using DeCon-ML-L and the original SlotCon. A consistent improvement is seen with our encoder-decoder framework, confirming the robustness of our method to more modern and bigger backbones. Notably, the performance of the model as well as the performance improvement appear larger than for a smaller ResNet-50 backbone.
Table 3: Performance of DeCon-SL ($\alpha=0.5$) adaptation of DenseCL and PixPro using a ResNet-50 encoder. The architectures were pre-trained for 800 epochs on COCO, and then fine-tuned on Pascal VOC and Cityscapes semantic segmentation. All fine-tuning results were averaged over three runs.
@@
Generalizability across frameworks: Table 3 presents semantic segmentation fine-tuning results of models pre-trained with DenseCL and PixPro, and their DeCon-SL adaptation. The performance of the model pre-trained with a decoder is improved compared to the base framework and improves even further when the pre-trained decoder is transferred, supporting the hypothesis that DeCon could be adapted to multiple frameworks. Transferring the pre-trained decoder however, is not always beneficial to downstream performance in the case of SlotCon. More details can be found in Supplementary Material 5.
Out-of-domain tasks: Table 4 depicts the fine-tuning performance of a model pre-trained on COCO to out-of-domain datasets. We observed a consistent improvement of the DeCon-SL pre-training adaptation compared to the encoder-only pre-training for both REFUGE and ISIC; moreover, this gain increased when the dataset size decreased. Further improvement was achieved when transferring the pre-trained decoder along with the encoder for ISIC dataset. Additionally, pre-training only on ISIC dataset consistently resulted in better performance compared to pre-training on COCO dataset. However, pre-training on REFUGE dataset did not outperform COCO pre-trained approaches, potentially due to the lack of available images (only 400) to develop a strong enough representation during pre-training. In both cases, DeCon-SL pre-training adaptation outperformed encoder-only pre-training, even if only the encoder was transferred. We also observed that pre-trained models outperformed models trained with randomly initialized weights. This gap was more notable in REFUGE dataset (over 20%).
Table 4: Transfer performance on out-of-domain datasets using randomly initialized or pre-trained weights. We pre-trained SlotCon and its DeCon-SL adaptation with a ResNet-50 encoder and an FCN decoder ($\alpha=0.5$) on COCO, REFUGE and ISIC training datasets. We fine-tuned on REFUGE and ISIC datasets under various dataset sizes. We averaged the fine-tuning mIoU results over three runs.

### v3

@@ Abstract
Contrastive learning methods in self-supervised settings have primarily focused on pre-training encoders, while decoders are typically introduced and trained separately for downstream dense prediction tasks. However, this conventional approach overlooks the potential benefits of jointly pre-training both encoder and decoder. In this paper, we propose DeCon, an efficient encoder-decoder self-supervised learning (SSL) framework that supports joint contrastive pre-training. We first extend existing SSL architectures to accommodate diverse decoders and their corresponding contrastive losses. Then, we introduce a weighted encoder-decoder contrastive loss with non-competing objectives to enable the joint pre-training of encoder-decoder architectures. By adapting a contrastive SSL framework for dense prediction, DeCon establishes consistent state-of-the-art performance on most of the evaluated tasks when pre-trained on Imagenet-1K, COCO and COCO+. Notably, when pre-training a ResNet-50 encoder on COCO dataset, DeCon improves COCO object detection and instance segmentation compared to the baseline framework by +0.37 AP and +0.32 AP, respectively, and boosts semantic segmentation by +1.42 mIoU on Pascal VOC and by +0.50 mIoU on Cityscapes. These improvements generalize across recent backbones, decoders, datasets, and dense tasks beyond segmentation and object detection, and persist in out-of-domain scenarios, including limited-data settings, demonstrating that joint pre-training significantly enhances representation quality for dense prediction. Code is available at https://github.com/sebquetin/DeCon.git.
**footnotetext: Equal contributions. Co-first authors are ordered in alphabetical order. Corresponding author: T. Ghosh (tapotosh.ghosh@ucalgary.ca)
@@ 1 Introduction
The growing demand for efficient Deep Learning (DL) methods stems from their ability to address complex tasks. While traditional approaches like supervised learning can train high-performing models, they depend on large volumes of high-quality annotated data. Obtaining such annotations is often labor-intensive, expensive, and sometimes impractical, even when abundant data is available. A common workaround is to use pre-trained models, typically encoders trained from scratch on large-scale datasets, to boost performance on downstream tasks with limited labels. Typically, an encoder is pre-trained to capture a representation of the input data, which can then be fine-tuned for downstream tasks such as classification or segmentation. During the downstream fine-tuning phase, randomly initialized layers are typically added on top of the encoder, and the complete architecture is retrained using a supervised learning approach on an annotated dataset. ImageNet [18] is a widely utilized labeled dataset for such pre-training. However, models pre-trained on ImageNet classification task often transfer poorly to downstream dense prediction tasks, such as detection and segmentation [65, 53, 45].
@@
In this work, we propose a Decoder-aware contrastive learning (DeCon) approach, an encoder-decoder framework that jointly pre-trains encoder and decoder in an SSL manner. DeCon enhances the representation power of the encoder, and prepares it more effectively for integration with decoders in downstream dense prediction tasks. The contributions of this paper are as follows:
  • We propose DeCon-SL, a single-level joint encoder-decoder adaptation of contrastive dense SSL frameworks. We show that jointly pre-training the encoder and decoder enhances the representation power of the encoder and improves downstream dense prediction performance, even when transferring only the pre-trained encoder.
  • We further extend this adaptation by introducing DeCon-ML, which builds on DeCon-SL with a multi-level decoder loss function and an encoder-level channel dropout to promote a comprehensive encoder feature utilization.
  • DeCon achieves new state-of-the-art (SOTA) results in COCO object detection and instance segmentation, and Pascal VOC, Cityscapes and ADE20K semantic segmentation when pre-trained on ImageNet-1K and COCO+.
  • We show that DeCon consistently improves performance across different backbones, a range of dense downstream tasks beyond object detection, semantic and instance segmentation as well as in various out-of-domain scenarios.
  • We show that DeCon can achieve gains without increasing parameter count and with comparable GPU cost relative to the original framework, e.g. SlotCon [55].
2 Related Work
@@
Recent SSL methods prioritize generalizability and move away from hand-crafted pretext tasks. These approaches can be broadly categorized into generative and contrastive methods. Generative approaches learn representations by modeling the underlying data distribution. This is typically achieved through reconstruction-based methods, where the model predicts corrupted parts of the input [24, 1, 20, 61], or through adversarial methods, where a generator model learns to synthesize realistic samples to challenge a discriminator, thereby encouraging the encoder to extract informative semantic features [57, 29]. Contrastive methods, which are the focus of this research, learn data representations by maximizing the similarity between features extracted from two augmented views of the same image. To avoid model collapse—a phenomenon where the model maps all inputs to the same representation—strategies such as negative sampling [25, 10], stop-gradient mechanisms [11], variance regularization [3, 4], momentum encoders [25], and other architectural and optimization techniques have been employed [6].
@@
Encoder-only versus encoder-decoder pre-training: Encoder-decoder architectures with skip connections between the encoder and decoder layers (i.e., the U-Net family of models) have achieved state-of-the-art results in many segmentation tasks [48, 30, 64]. A common approach for pre-training such models in a self-supervised manner is image reconstruction [33, 5, 8]. However, the presence of skip connections allows a substantial amount of information to bypass the encoder’s final layer—often referred to as the bottleneck layer–resulting in lower-quality embeddings generated by the encoder. Our proposed approach preserves skip connections while addressing the issue of inferior representations generated by the decoder.
Encoder-only contrastive learning frameworks are built solely around an encoder, incorporating components such as predictors and projectors. However, they do not include a decoder. After pre-training, a randomly initialized decoder is attached to the pre-trained encoder in dense prediction downstream tasks to form a complete architecture. While some generative SSL frameworks use a decoder for pre-training [17, 43, 24], a clear gap remains in exploring vision model pre-training that jointly employs encoder and decoder losses within a unified contrastive learning framework to learn informative data representations.
@@ 3.1 DeCon-SL
Given a vision SSL framework that pre-trains an encoder using a teacher-student architecture [28], we retain the encoders and any existing “auxiliary layers” (e.g., predictors, projection heads), and add decoders and their corresponding auxiliary layers for both teacher and student networks. For all models, the decoder auxiliary layers are designed to match those of the encoder, with adjustments applied to align with the decoder’s output feature size. We obtain two losses, one is computed from the encoder features while the other one is computed from the decoder features. Consequently, the loss function is defined as follows:
@@ 3.3 Implementation details
Architecture: We use a ResNet-50 encoder [26] as a backbone and implement a Fully Convolutional Network (FCN) [37] and a skip-connected Feature Pyramid Network (FPN) [34] as two decoder architectures in the DeCon framework. We follow the mmseg [14] FCN implementation from SlotCon [55] and customize the Detectron2 [56] FPN implementation. Details on these decoders are provided in Supplementary Material Section 1. We use DeCon to adapt the SlotCon [55] framework. Additional experiments are conducted using DenseCL [53] and PixPro [60] framework, or a ConvNeXt-Small [36] backbone.
@@
We adapt in a similar way DenseCL and PixPro into a DeCon-SL framework using an FCN decoder. More details are given in Supplementary Material Section 2. We adapt DenseCL implementation from mmselfsup package [15] version 1.0.0 and PixPro from the original implementation.
@@
Although the pre-trained backbone remains unchanged, adding multiple decoder levels and associated loss functions increases the overall number of parameters in the DeCon-ML pre-training. To enable a fair comparison with the original SlotCon framework using a similar parameter budget, we implemented a reduced version of DeCon-ML that includes only the first two decoder levels originating from the encoder bottleneck. Additionally, the hidden dimension in the decoder projector was reduced from 4096 to 2048. This smaller version, which matches the parameter count of the original SlotCon framework (when $\alpha=0$), is referred to as DeCon-ML-S, while the full model with four decoder levels and unmodified configuration is denoted DeCon-ML-L. Detailed figures of DeCon-SL, DeCon-ML-S and DeCon-ML-L adaptations of SlotCon are available in the Supplementary Material Section 3.
@@
Pre-training setup: We pre-train DeCon on COCO 2017 dataset [35], COCO+ (train2017 + unlabeled2017 datasets) and ImageNet-1K datasets [18]. The augmentations and hyperparameters—learning rate (LR), weight decay, schedulers, input shape, momentum, and so on—are kept the same as the adapted SSL framework. Models are pre-trained on COCO and COCO+ for 800 epochs, and for 200 epochs on ImageNet-1K. The SlotCon ResNet-50 models are pre-trained using a LARS optimizer, with a total batch size of 512 split across all available GPUs. The base LR is 1.0 scaled with the batch size following [55] and updated with a cosine LR decay schedule using a weight decay of $10^{-5}$ and 5 warm-up epochs. The ConvNeXt-Small backbone is pre-trained on COCO dataset with SlotCon and DeCon-ML using an AdamW optimizer [38], a base LR of 0.002 and weight-decay 0.05. To compare with ViT-based methods, we also pre-train a ConvNeXt-S backbone on ImageNet-1K for 250 epochs with DeCon-SL framework and an FCN decoder using a base LR of 0.001 and drop path rate of 0.1. Remaining hyperparameters are kept the same as for the ResNet-50 implementation. The pre-trainings using ConvNeXt-S backbones or FPN decoders are done using eight 80GB NVIDIA H100 GPUs. The pre-trainings without decoder and with FCN decoders are conducted on one 80GB NVIDIA A100 GPU.
To evaluate the generalization of our approach to different SSL frameworks, we adapt DenseCL and PixPro frameworks with DeCon-SL by pre-training a ResNet-50 encoder with an FCN decoder on COCO 2017 dataset. Optimization follows the base frameworks. To evaluate the dependence to the pre-training dataset, we also pre-train a similar DeCon-SL adaptation of SlotCon on REFUGE [41] and ISIC [13] datasets. We provide all pre-training details in Supplementary Material Section 2.
@@ 3.4 Evaluation protocol
We evaluate the different pre-trained models on various downstream tasks. Unless stated otherwise, we only transfer the pre-trained encoder for the fine-tuning experiments. We initialize the network with a pre-trained encoder and fine-tune end-to-end in all cases. All fine-tuning results reported are averaged over three independent runs. The pre-trainings, however, are only run once. All fine-tuning experiments related to the same downstream task are run on the same hardware. We perform semantic segmentation tasks using mmsegmentation [14] version 0.30.0. and other remaining tasks using Detectron2 package [56] version 0.6.
Object detection and instance segmentation: We fine-tune a Mask R-CNN with a pre-trained ResNet-50 or ConvNeXt-Small backbone on COCO 2017 dataset [35].
Semantic segmentation: We fine-tune a pre-trained ResNet-50 or ConvNeXt-Small with an FCN decoder on Pascal VOC [21], and ResNet-50 with an FCN decoder on ADE20K [63] and Cityscapes [16] following SlotCon [55] and PixCon [42]. We also fine-tune an ImageNet-1K-pre-trained ConvNeXt-Small backbone with an UPerNet decoder on ADE20K dataset.
Panoptic segmentation, keypoint detection, and dense pose estimation: We adapt task-specific Detectron2 configurations based on the SlotCon object detection architecture, and fine-tune on the COCO 2017 dataset for panoptic segmentation and keypoint detection, and on the COCO 2014 dataset for dense pose estimation.
Generalization to out-of-domain datasets: We fine-tune a ResNet-50-FCN architecture on REFUGE [41] and ISIC 2017 [13] datasets for medical semantic segmentation using 5, 25, and 100% of the training data. We also fine-tune a ResNet-50-deeplabv3+ architecture for PlantSeg  [54] semantic segmentation and a Faster R-CNN ResNet-50-FPN architecture for object detection on PlantDoc [49] and Detecting Disease  [47] datasets. Fine-tuning on agriculture datasets is performed using 10 and 100% of the datasets.
More details on the training hyperparameters and target datasets are provided in Supplementary Material Section 2.
@@ 4.1 Experiments
Table 1: Performance of DeCon-SL ($\alpha=0.25$) and DeCon-ML (Small/Large, $\alpha=0$, dropout=0.5) adaptations of SlotCon with a ResNet-50 backbone pre-trained for 800 epochs on COCO/COCO+ and 200 epochs on ImageNet-1K, then fine-tuned on different downstream tasks. Only pre-trained encoders were transferred. PixCon ImageNet-1K results are not reported due to unavailable checkpoints. Results are averaged over three runs. Re-implemented, * Results from original paper, $\lozenge$ Result from SlotCon [55] paper.
Pret. Dataset Framework Backbone Pret. Dec. Object Det. Instance Seg. Semantic Seg. Dense pose Est. Panoptic Seg. Keypoint Det.
COCO COCO VOC City ADE COCO COCO COCO
AP AP50 AP75 AP AP50 AP75 mIoU mIoU mIoU GPS AP GPSm AP PQ AP
[…]
COCO PixCon (2024) [42]* ResNet-50 - 40.81 60.97 44.80 36.80 57.93 39.62 72.95 76.62 38.00 - - - -
SoCo-D (2024) [44]* - 40.3 60.1 44.0 35.1 56.9 37.6 - - - - - - -
SlotCon (2022) [55] - 40.81 60.95 44.37 36.80 57.98 39.54 71.50 75.95 38.57 63.39 64.84 40.31 65.66
DeCon-SL FCN 40.97 61.22 44.81 36.92 58.12 39.78 73.01 76.21 38.81 63.50 64.98 40.44 65.77
DeCon-ML-S FPN 40.97 61.20 44.71 36.94 58.20 39.63 72.80 76.21 38.36 63.75 65.08 40.52 65.66
DeCon-ML-L FPN 41.18 61.38 44.91 37.12 58.35 39.94 72.92 76.45 38.70 63.72 64.94 40.90 65.88
COCO+ PixCon (2024) [42]* ResNet-50 - 41.2 - - 37.1 - - 73.9 77.00 38.80 - - - -
SlotCon (2022) [55] - 41.63 62.10 45.67 37.57 59.07 40.45 73.93 76.43 39.11 - - - -
DeCon-SL FCN 41.86 62.43 45.73 37.75 59.40 40.48 74.46 76.65 39.25 - - - -
DeCon-ML-L FPN 42.08 62.42 46.13 37.84 59.41 40.75 75.36 77.00 39.04 - - - -
ImageNet-1K Supervised $\lozenge$ ResNet-50 - 39.7 59.5 43.3 35.9 56.6 38.6 74.4 74.6 37.9 - - - -
DINO (2021) [6] - 40.24 60.25 44.13 36.47 57.49 39.20 73.09 75.57 37.30 - - - -
SlotCon (2022) [55] - 41.69 62.07 45.59 37.59 58.97 40.49 75.02 76.15 38.97 - - - -
DeCon-ML-L FPN 41.80 62.12 45.73 37.73 59.08 40.68 75.40 76.51 39.01 - - - -
Table 2: Performance of DeCon-ML-L ($\alpha=0$, dropout = 0.5) and SlotCon with ConvNeXt-S encoder, pre-trained for 800 epochs on COCO and fine-tuned on segmentation and detection tasks.
@@
Table 3: Comparison of ViT and ConvNeXt models pre-trained on ImageNet-1K. We pre-trained a ConvNeXt-S encoder using DeCon-SL (FCN decoder, $\alpha=0.25$). All models are fine-tuned with an UPerNet decoder [58] on ADE20K. *Results from MixedAE [7]; †Results using two decoder levels to match with DeCon-SL decoder.
Framework Pret. Backb. Backb. Params. (M) Pret. Epoch ADE20k (mIoU)
MoCov3 (2021) [12]* ViT-B 88 600 46.8
DINO (2021) [6]* ViT-B 88 1600 46.9
BEiT (2021) [2]* ViT-B 88 300 44.7
MAE (2022) [24]* ViT-B 88 300 46.7
MixedAE (2023) [7]*† ViT-B 88 300 47.4
DeCon-SL [ours] ConvNeXt-S 50 250 48.02
ResNet-50 backbone: Table 1 and Supplementary Material Table S1 present the results of fine-tuning a ResNet-50 backbone pre-trained with various contrastive SSL frameworks. We mainly compare with SlotCon and PixCon as they demonstrated the most competitive performance in dense downstream tasks among previously published contrastive methods. DeCon-ML-L is consistently establishing a new SOTA for all tasks when pre-trained on ImageNet-1K and is consistently improving upon its base framework SlotCon and showing SOTA results for most of the evaluated downstream tasks when pre-trained on COCO and COCO+ dataset. Notably, it reduces the gap between PixCon and SlotCon for Pascal VOC and Cityscapes semantic segmentation when pre-trained on COCO. This supports the fact that the DeCon adaptation is robust across pre-training datasets and fine-tuning tasks. Moreover, DeCon-ML-S and DeCon-SL, whose training costs are comparable to SlotCon exhibit better performance than SlotCon and other contrastive frameworks. Finally, Tab. 1 also shows that performance improvement compared to SlotCon are robust to other downstream tasks such as dense pose estimation, panoptic segmentation and keypoint detection.
ConvNeXt backbone and comparisons with ViT: Table 2 shows the fine-tuning performance of a ConvNeXt-Small backbone pre-trained using DeCon-ML-L and the base SlotCon framework. A consistent improvement is seen with our encoder-decoder framework, confirming the robustness of our method to more modern and bigger backbones. Notably, both the absolute performance and the relative improvement are greater than those observed with the ResNet-50 backbone. Table 3 presents a performance comparison between a ConvNeXt-S model pre-trained on ImageNet-1K with DeCon-SL and ViT backbones pre-trained using various frameworks. Despite its smaller size and shorter pre-training schedule, the ConvNeXt-S backbone, pre-trained for dense tasks, outperforms all competing methods.
Table 4: Performance of DeCon-SL ($\alpha=0.5$) adapting DenseCL and PixPro with a ResNet-50 encoder, pre-trained for 800 epochs on COCO and fine-tuned on Pascal VOC and Cityscapes.
@@
Generalizability across frameworks: Table 4 presents semantic segmentation fine-tuning results for models pre-trained with DenseCL and PixPro, along with their DeCon-SL adaptations. Pre-training with a decoder improves performance over the base frameworks, and transferring the decoder further enhances results, supporting the hypothesis that DeCon can be adapted to multiple frameworks. However, transferring the pre-trained decoder does not consistently benefit downstream performance in the case of SlotCon, as seen in Tab. 5 and Supplementary Material 5.
Out-of-domain tasks: Table 5 shows fine-tuning results of COCO-pre-trained encoders on REFUGE and ISIC. DeCon-SL pre-training consistently outperforms SlotCon, with larger gains under limited data. Transferring both encoder and decoder further improves performance on ISIC. Additionally, pre-training only on ISIC dataset consistently resulted in better performance compared to pre-training on COCO dataset. However, pre-training on REFUGE dataset did not outperform COCO pre-trained approaches, potentially due to the lack of available images (only 400) to develop a strong enough representation during pre-training. In both ISIC and REFUGE pre-trainings, DeCon-SL outperformed SlotCon. Table 6 shows that DeCon variants also perform better than SlotCon in agricultural object detection and semantic segmentation tasks using 10 or 100% of the datasets. DeCon pre-trained models also outperformed randomly initialized ones.
Table 5: Transfer performance with SlotCon and DeCon-SL (ResNet-50, FCN, $\alpha{=}0.5$) pre-trained on COCO, REFUGE, and ISIC, then fine-tuned on out-of-domain medical datasets (REFUGE and ISIC sem. seg. ) in various data settings.
@@
mIoU mIoU mIoU mIoU mIoU mIoU mIoU mIoU